# Towards Domain Specification of Embedding Models in Medicine

### v1

@@ Abstract
{IEEEkeywords}
Embedding, Benchmark, Medicine, Biology, Evaluation
footnotetext: 1Engineering Technology Building, McMaster Universityfootnotetext: These authors contributed equally to this workfootnotetext: Corresponding author: Mohammad Khodadad, khodam3@mcmaster.ca
@@ 2 Literature review
Recent work merges unsupervised and supervised contrastive objectives. GTE [16] uses large‐scale unsupervised contrastive pretraining followed by supervised fine‐tuning. Nomic-Embed [17] combines masked language modeling and contrastive learning over contexts up to 8,192 tokens with FlashAttention, gradient caching, and DeepSpeed. These steps shift from bidirectional pretraining toward specialized contrastive methods, motivating our medical embedding benchmark.
@@
Despite rapid advances in general text embeddings, evaluations of medical‐domain models remain fragmented and narrow. Most benchmarks cover only a few clinical datasets or task types, leaving critical gaps in our understanding of model behavior across real‐world healthcare scenarios. The Massive Text Embedding Benchmark (MTEB) [8] addresses this by unifying 58 datasets in 112 languages over eight core tasks, including bitext mining, classification, clustering, pair classification, reranking, retrieval, semantic textual similarity, and summarization via an open-source design and Hugging Face leaderboard. However, MTEB’s medical subsection comprises only 12 unevenly sized datasets (eight in English), limiting its utility for rigorous medical embedding assessment. The Chemical Text Embedding Benchmark (ChemTEB) [27] demonstrates how specialized suites can capture domain‐specific challenges through tasks like chemical document classification, retrieval, and bitext mining, inspiring our approach to medical text embedding evaluation.
Subsequent efforts have targeted scalable and retrieval‐focused evaluation in clinical contexts. Soffer et al. [28] tested 30 top MTEB models on heterogeneous corpora—PubMed abstracts, LLaMA‐3‐70B–generated synthetic EHRs, Mount Sinai patient records, and MIMIC-IV notes—distinguishing “Short Tasks” (e.g., triage notes) from “Long Tasks” (e.g., full patient histories). While revealing length‐dependent performance trends, this framework remains closed‐source and limited to matching tasks. HEmTEB (Health Embedding Benchmark) [29] refines retrieval evaluation by curating clinician‐supervised datasets across ten specialties and assessing cross‐lingual capabilities, yet it too omits classification, clustering, and semantic similarity assessments. Parallel evaluations on specialized tasks (e.g., biomedical terminology mapping) similarly find that no single model excels uniformly across all semantic alignment challenges, underscoring the need for broader benchmarks [30].
These specialized frameworks underscore the value of targeted benchmarks but also highlight the need for a more expansive, standardized medical embedding evaluation. We therefore propose the Medical Text Embedding Benchmark—a large‐scale, open‐source suite spanning classification, clustering, retrieval, semantic similarity, and beyond, drawn from diverse clinical and biomedical corpora—to provide a comprehensive, transparent, and actionable performance assessment for medical NLP models.
@@ 3.1 Sources of data for MedTEB and MedTE
Our corpus integrates complementary strands of biomedical knowledge from peer-reviewed literature, preprint archives, real-world clinical narratives, structured registries, human-curated encyclopedic resources, standardized question–answer benchmarks, and specialized training corpora. We begin with PubMed [31], a comprehensive catalogue of biomedical abstracts, and its full-text counterpart PubMed Central (PMC) [32], anchoring our collection in rigorously vetted scientific discourse. To capture authentic clinical language we incorporate MIMIC-IV [33], a de-identified EHR dataset rich in clinical notes, discharge summaries, and structured patient data reflective of bedside documentation. Prospective study descriptions from ClinicalTrials.gov [34] contribute structured trial metadata—objectives, interventions, and eligibility criteria—extending our coverage to both ongoing and completed research across therapeutic domains. For rapid insight into emerging findings we harvest life-science preprints from bioRxiv [35] and clinically oriented preprints from medRxiv [36], whose detailed metadata support fine-grained filtering and trend analysis.
@@ 3.2 MedTE Training
For every corpus, we derive semantically aligned positive pairs—two text fragments that convey the same underlying meaning. The pairing heuristics for each data source are listed in Table 1. All texts are first lower‑cased and then tokenised with the GTE tokenizer (a bert‑base‑uncased variant) to ensure a uniform sub‑word vocabulary.
@@
Table 1: Positive-pair construction for self-supervised contrastive learning (numbers rounded).
@@
where $z=\operatorname{MeanPool}(\text{GTE}(x))$ is the sentence embedding, $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a learned temperature parameter. No manually labelled negatives are required: all non‑matching sentences in the same batch act as implicit negatives, enabling scalable self‑supervision. To keep domain shift under control, the data loader assembles batches from a single corpus at a time, so every negative originates from the same data source as its corresponding positive pair.
@@
Training is performed on 2,033,800 positive pairs (1) drawn proportionally from all corpora. Each optimization step processes a batch of 1024 pairs, and we train for 8000 steps, which we found sufficient for convergence. We use the AdamW [42] optimizer with weight decay of 0.010.010.01; the learning rate is linearly warmed up to its peak over the first 1,000 steps and then decays according to a cosine schedule. Mixed precision (bfloat16) is enabled to reduce memory footprint, and gradient checkpointing is activated to fit larger batches without compromising sequence length. We also leverage DeepSpeed[43] for efficient large‐batch training and accelerated throughput. At inference time, embeddings are $L_{2}$-normalized so that inner-product search is equivalent to cosine similarity.
@@ 3.3 MedTEB Development
In our benchmark, we assembled 51 datasets across four categories—Classification, Clustering, Pair Classification, and Retrieval—to comprehensively evaluate the representational power of embedding models. For Classification, given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune the embedding model on the training split $\mathcal{D}_{\mathrm{train}}$, feed the resulting sentence embeddings into a logistic regression classifier trained on the same split, and report macro-averaged $F_{1}$ on the test split $\mathcal{D}_{\mathrm{test}}$. In Clustering tasks, each input $x_{i}$ is embedded into $z_{i}\in\mathbb{R}^{d}$ and Mini-Batch $k$-means (batch size 32) is applied to produce cluster labels $\hat{c}_{i}$, with clustering quality measured against true labels $c_{i}$ using the V-measure. For Pair Classification, given $\mathcal{D}=\{(x_{i}^{(1)},x_{i}^{(2)},y_{i})\}_{i=1}^{N}$ with binary labels $y_{i}$, we compute four similarity or distance scores (cosine, Euclidean, Manhattan, and dot product) per pair, select a threshold $\tau$ on the training data to maximize $F_{1}$, and define predictions $\hat{y}_{i}$ accordingly, reporting the best $F_{1}$ across all metrics. Finally, Retrieval tasks consist of query sets $\mathcal{Q}$, document corpora $\mathcal{D}$, and relevance pairs $\mathcal{R}\subseteq\mathcal{Q}\times\mathcal{D}$; we embed every query and document into vectors $z_{q_{j}}$ and $z_{d_{k}}$, rank documents by descending cosine similarity $\cos(z_{q_{j}},z_{d_{k}})$, and measure performance with nDCG@10 averaged over queries. This design enables a holistic analysis of embedding quality across supervised and unsupervised scenarios, capturing semantic grouping, pairwise similarity discrimination, and practical retrieval capabilities in information-seeking tasks. Table 2 provides an overview of all datasets and their properties.
Task MIMIC-IV PMC PubMed Wikipedia MedQA MedMCQA MedQUAD medRxiv bioRxiv Total
Classification 15
Clustering 12
Pair Classification 12
Retrieval 12
@@
[…]
@@
[…]
We evaluated a diverse collection of embedding models that differ along two key dimensions—domain (general versus medical) and training objective (contrastive learning versus standard pretraining). Table 3 summarizes each model’s size, parameter count, maximum context length, embedding dimensionality, and whether it was trained with a contrastive objective. The non-medical group includes popular transformers such as BGE, SciBERT, and MiniLM variants, while the medical group comprises specialized models like ClinicalBERT, BioSimCSE, and our own MedTE.
Model Name Size (MB) Params (M) Context Emb. Dim Cons. Lear.
Non‑medical models
BAAI BGE Base En V1.5 418 109.48 512 768 Yes
AllenAI SciBERT Scivocab Uncased 422 109.92 512 768 No
Google BERT Base Uncased 420 109.48 512 768 No
Intfloat E5 Base 418 109.48 512 768 Yes
Nomic AI Nomic Embed Text V1 522 136.73 819 768 Yes
Nomic AI Nomic Embed Text V1 Unsupervised 522 136.73 512 768 Yes
Sentence‑Tfrs All‑MiniLM‑L6‑v2 87 22.71 512 384 Yes
Sentence‑Tfrs All‑MPNet‑Base‑v2 418 109.49 512 768 Yes
Thenlper GTE Base 209 109.48 512 768 Yes
Medical models
BioNLP BlueBERT PubMed MIMIC Uncased L‑12 H‑768 A‑12 420 109.48 512 768 No
Abhinand MedEmbed Base 420 109.48 512 768 Yes
Emily Alsentzer Bio‑ClinicalBERT 416 108.31 512 768 No
Kamalkraj BioSimCSE BioLinkBERT Base 413 108.23 512 768 Yes
Malteos SciNCL 419 109.92 512 768 Yes
MedicalAI ClinicalBERT 517 134.73 512 768 No
Microsoft BiomedBERT Base Uncased Abs+Fulltext 420 109.48 512 768 No
MedTE Cl15 Step 8000 438 109.48 512 768 Yes
Table 3: Specifications for Evaluated Embedding Models on MedTEB, with Domain and Contrastive‑Learning Indicators
@@ 4.1 Overall performance
Averaged across all 16 task metrics, Our model attains a mean score of 0.5780.5780.578, eclipsing the nearest competitor at 0.5250.5250.525 and establishing a new state of the art for medical sentence representation. These results underscore that in specialized domains such as healthcare, self-supervised contrastive learning is not optional but decisive, future medical NLP systems should therefore prioritize contrastive frameworks to achieve consistent improvements in classification, clustering, and retrieval.
@@ 4.2 Classification
Our model achieves an F1 score of 0.72±0.230.72$\pm$ 0.230.72 ± 0.23, outperforming both the strongest non-medical baseline (MPNet Base V2, 0.70±0.220.70$\pm$ 0.220.70 ± 0.22) and the leading clinical encoder (Abhinand MedEmbed Base, 0.69±0.220.69$\pm$ 0.220.69 ± 0.22). In practical terms, this +0.02+0.02+ 0.02+0.03+0.03+ 0.03 margin translates to hundreds fewer misclassifications in downstream triage workflows, since each point of F1 corresponds to hundreds of patient records.
@@ 4.3 Clustering
Using the V-measure, our model records 0.38±0.240.38$\pm$ 0.240.38 ± 0.24, a gain of +0.04+0.04+ 0.04 over the top non-medical contrastive model (Nomic Embed Text V1 Unsupervised, 0.34±0.260.34$\pm$ 0.260.34 ± 0.26) and +0.02+0.02+ 0.02 above the best domain-tuned rival (SciNCL, 0.36±0.260.36$\pm$ 0.260.36 ± 0.26).
@@ 4.4 Pair Classification and Retrieval
For semantically equivalent sentence detection, posts an F1 of 0.74±0.170.74$\pm$ 0.170.74 ± 0.17, beating MPNet Base V2 by +0.07+0.07+ 0.07 and MedEmbed Base by +0.06+0.06+ 0.06. In retrieval, it achieves nDCG@10 =0.45±0.32=0.45$\pm$ 0.32= 0.45 ± 0.32, improving over MPNet Base V2 by +0.10+0.10+ 0.10 and the strongest clinical baseline by +0.06+0.06+ 0.06. These gains reflect a denser, better-calibrated embedding space that surfaces relevant clinical evidence with fewer query reformulations.
@@ 5.1 Per Source Performance on MedTE
As explained in the data section, we created our benchmark using a diverse set of medical text data from different sources. In table 5, we visualized the average performance of models on different sources of data. Both Abhinand MedEmbed Base and BAAI Bge Base En V1.5 achieve particularly strong performance on PubMed (0.89–0.90) and BioRxiv (0.89–0.90), substantially outpacing models such as SciBERT (PubMed 0.50; BioRxiv 0.18) and BioClinicalBERT (PubMed 0.47; BioRxiv 0.10). Ours further refines this trend, achieving the highest MIMIC‐IV score (0.61) and maintaining robust performance across other clinical sources (Clinical Trials 0.81; MedRxiv 0.74). In comparison, general embeddings such as All MiniLM L6 V2 and MPNet Base V2 hover around 0.51–0.55 on PubMed and just 0.64–0.67 on BioRxiv, underscoring the importance of specialized pretraining for domain‐specific tasks.
The MedMCQA and MedQA tasks involve multi‐option question answering, which poses a significant challenge for embedding models to capture the relationship between a question and its correct answer. All models converge to a 0.39 average on these datasets, highlighting this difficulty. Nevertheless, Ours outperforms competitors on MedMCQA, likely attributable to its extensive pretraining on a diverse set of medical data. On the general‐domain Wikipedia task, domain‐adapted models incur only a minor drop—Ours achieves 0.45 and Thenlper GTE Base reaches 0.43—outperforming many general embeddings (All MiniLM L6 V2 at 0.41; BERT Base at 0.26). Overall, these results confirm that our model outperforms existing embeddings across specialized biomedical evaluations while preserving—and in some cases enhancing—performance on broader‐scope information sources.
@@
Figure 1: Comparison of model performance on Wikipedia versus MIMIC‑IV

### v2

@@ Abstract
Keywords Embedding, Benchmark, Medicine, Biology, Evaluation
@@ 2 Literature review
Recent work merges unsupervised and supervised contrastive objectives. GTE [16] uses large-scale unsupervised contrastive pretraining followed by supervised fine-tuning. Nomic-Embed [17] combines masked language modeling and contrastive learning over contexts up to 8,192 tokens with FlashAttention, gradient caching, and DeepSpeed. These steps shift from bidirectional pretraining toward specialized contrastive methods, motivating our medical embedding benchmark.
@@
Despite rapid advances in general text embeddings, evaluations of medical-domain models remain fragmented and narrow. Most benchmarks cover only a few clinical datasets or task types, leaving critical gaps in our understanding of model behavior across real-world healthcare scenarios. The Massive Text Embedding Benchmark (MTEB) [8] addresses this by unifying 58 datasets in 112 languages over eight core tasks, including bitext mining, classification, clustering, pair classification, reranking, retrieval, semantic textual similarity, and summarization via an open-source design and Hugging Face leaderboard. However, MTEB’s medical subsection comprises only 12 unevenly sized datasets (eight in English), limiting its utility for rigorous medical embedding assessment. The Chemical Text Embedding Benchmark (ChemTEB) [27] demonstrates how specialized suites can capture domain-specific challenges through tasks like chemical document classification, retrieval, and bitext mining, inspiring our approach to medical text embedding evaluation.
Subsequent efforts have targeted scalable and retrieval-focused evaluation in clinical contexts. Soffer et al. [28] tested 30 top MTEB models on heterogeneous corpora, PubMed abstracts, LLaMA-3-70B–generated synthetic EHRs, Mount Sinai patient records, and MIMIC-IV notes, distinguishing “Short Tasks” (e.g., triage notes) from “Long Tasks” (e.g., full patient histories). While revealing length-dependent performance trends, this framework remains closed-source and limited to matching tasks. HEmTEB (Health Embedding Benchmark) [29] refines retrieval evaluation by curating clinician-supervised datasets across ten specialties and assessing cross-lingual capabilities, yet it too omits classification, clustering, and semantic similarity assessments. Parallel evaluations on specialized tasks (e.g., biomedical terminology mapping) similarly find that no single model excels uniformly across all semantic alignment challenges, underscoring the need for broader benchmarks [30].
These specialized frameworks underscore the value of targeted benchmarks but also highlight the need for a more expansive, standardized medical embedding evaluation. We therefore propose the Medical Text Embedding Benchmark, a large-scale, open-source suite spanning classification, clustering, retrieval, semantic similarity, and beyond, drawn from diverse clinical and biomedical corpora, to provide a comprehensive, transparent, and actionable performance assessment for medical NLP models.
@@ 3.1 Sources of data for MedTEB and MedTE
Our corpus integrates complementary strands of biomedical knowledge from peer-reviewed literature, preprint archives, real-world clinical narratives, structured registries, human-curated encyclopedic resources, standardized question–answer benchmarks, and specialized training corpora. We begin with PubMed [31], a comprehensive catalogue of biomedical abstracts, and its full-text counterpart PubMed Central (PMC) [32], anchoring our collection in rigorously vetted scientific discourse. To capture authentic clinical language we incorporate MIMIC-IV [33], a de-identified EHR dataset rich in clinical notes, discharge summaries, and structured patient data reflective of bedside documentation. Prospective study descriptions from ClinicalTrials.gov [34] contribute structured trial metadata, objectives, interventions, and eligibility criteria, extending our coverage to both ongoing and completed research across therapeutic domains. For rapid insight into emerging findings we harvest life-science preprints from bioRxiv [35] and clinically oriented preprints from medRxiv [36], whose detailed metadata support fine-grained filtering and trend analysis.
@@ 3.2 MedTE Training
For every corpus, we derive semantically aligned positive pairs, two text fragments that convey the same underlying meaning. The pairing heuristics for each data source are listed in Table 1. All texts are first lower-cased and then tokenised with the GTE tokenizer (a bert-base-uncased variant) to ensure a uniform sub-word vocabulary.
@@
Table 1: Positive-pair construction for MedTE’s unsupervised contrastive learning (numbers rounded).
@@
where $z=\operatorname{MeanPool}(\text{GTE}(x))$ is the sentence embedding, $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a learned temperature parameter. No manually labelled negatives are required: all non-matching sentences in the same batch act as implicit negatives, enabling scalable self-supervision. To keep domain shift under control, the data loader assembles batches from a single corpus at a time, so every negative originates from the same data source as its corresponding positive pair.
@@
Training is performed on 2,033,800 positive pairs (1) drawn proportionally from all corpora. Each optimization step processes a batch of 1024 pairs, and we train for 8000 steps, which we found sufficient for convergence. We use the AdamW [42] optimizer with weight decay of $0.01$; the learning rate is linearly warmed up to its peak over the first 1,000 steps and then decays according to a cosine schedule. Mixed precision (bfloat16) is enabled to reduce memory footprint, and gradient checkpointing is activated to fit larger batches without compromising sequence length. We also leverage DeepSpeed[43] for efficient large-batch training and accelerated throughput. At inference time, embeddings are $L_{2}$-normalized so that inner-product search is equivalent to cosine similarity.
@@ 3.3 MedTEB Development
In our benchmark, we assembled 51 datasets across four categories, Classification, Clustering, Pair Classification, and Retrieval, to comprehensively evaluate the representational power of embedding models. For Classification, given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune the embedding model on the training split $\mathcal{D}_{\mathrm{train}}$, feed the resulting sentence embeddings into a logistic regression classifier trained on the same split, and report macro-averaged $F_{1}$ on the test split $\mathcal{D}_{\mathrm{test}}$. In Clustering tasks, each input $x_{i}$ is embedded into $z_{i}\in\mathbb{R}^{d}$ and Mini-Batch $k$-means (batch size 32) is applied to produce cluster labels $\hat{c}_{i}$, with clustering quality measured against true labels $c_{i}$ using the V-measure. For Pair Classification, given $\mathcal{D}=\{(x_{i}^{(1)},x_{i}^{(2)},y_{i})\}_{i=1}^{N}$ with binary labels $y_{i}$, we compute four similarity or distance scores (cosine, Euclidean, Manhattan, and dot product) per pair, select a threshold $\tau$ on the training data to maximize $F_{1}$, and define predictions $\hat{y}_{i}$ accordingly, reporting the best $F_{1}$ across all metrics. Finally, Retrieval tasks consist of query sets $\mathcal{Q}$, document corpora $\mathcal{D}$, and relevance pairs $\mathcal{R}\subseteq\mathcal{Q}\times\mathcal{D}$; we embed every query and document into vectors $z_{q_{j}}$ and $z_{d_{k}}$, rank documents by descending cosine similarity $\cos(z_{q_{j}},z_{d_{k}})$, and measure performance with nDCG@10 averaged over queries. This design enables a holistic analysis of embedding quality across supervised and unsupervised scenarios, capturing semantic grouping, pairwise similarity discrimination, and practical retrieval capabilities in information-seeking tasks. Table 2 provides an overview of all datasets and their properties.
Task MIMIC-IV PMC PubMed Wikipedia MedQA MedMCQA MedQUAD medRxiv bioRxiv Total Classification 15 Clustering 12 Pair Classification 12 Retrieval 12
@@
For each dataset, we extracted the introductory or abstract portion of every document as the input text. Labels were then assigned by one of three methods: (i) the author’s own keywords, (ii) the document’s parent category in its source hierarchy, or (iii) the search term used to retrieve the document when no explicit label was available. This process yields a clean set of $\langle\text{text},\text{label}\rangle$ pairs that serve both classification and clustering experiments.
@@
We reused the same corpus by treating each original sentence as a “document” and its GPT-4o paraphrase (or, in QA datasets, the question) as a query. This guarantees that every $\langle\text{query},\text{document}\rangle$ pair is relevant, while the rest of the corpus acts as distractors. By mirroring the Pair Classification setup, we ensure consistent difficulty across both tasks.
We evaluated a diverse collection of embedding models that differ along two key dimensions, domain (general versus medical) and training objective (contrastive learning versus standard pretraining). Table 3 summarizes each model’s size, parameter count, maximum context length, embedding dimensionality, and whether it was trained with a contrastive objective. The non-medical group includes popular transformers such as BGE, SciBERT, and MiniLM variants, while the medical group comprises specialized models like ClinicalBERT, BioSimCSE, and our own MedTE.
Model Name Size (MB) Params (M) Context Emb. Dim Cons. Lear. Non-medical models BAAI BGE Base En V1.5 418 109.48 512 768 Yes AllenAI SciBERT Scivocab Uncased 422 109.92 512 768 No Google BERT Base Uncased 420 109.48 512 768 No Intfloat E5 Base 418 109.48 512 768 Yes Nomic AI Nomic Embed Text V1 522 136.73 819 768 Yes Nomic AI Nomic Embed Text V1 Unsupervised 522 136.73 512 768 Yes Sentence-Tfrs All-MiniLM-L6-v2 87 22.71 512 384 Yes Sentence-Tfrs All-MPNet-Base-v2 418 109.49 512 768 Yes Thenlper GTE Base 209 109.48 512 768 Yes Medical models BioNLP BlueBERT PubMed MIMIC Uncased L-12 H-768 A-12 420 109.48 512 768 No Abhinand MedEmbed Base 420 109.48 512 768 Yes Emily Alsentzer Bio-ClinicalBERT 416 108.31 512 768 No Kamalkraj BioSimCSE BioLinkBERT Base 413 108.23 512 768 Yes Malteos SciNCL 419 109.92 512 768 Yes MedicalAI ClinicalBERT 517 134.73 512 768 No Microsoft BiomedBERT Base Uncased Abs+Fulltext 420 109.48 512 768 No MedTE Cl15 Step 8000 438 109.48 512 768 Yes
Table 3: Specifications for Evaluated Embedding Models on MedTEB, with Domain and Contrastive-Learning Indicators
@@ 4.1 Overall performance
Averaged across all 16 task metrics, Our model attains a mean score of $0.578$, eclipsing the nearest competitor at $0.539$ (Abhinand MedEmbed Base) and establishing a new state of the art for medical sentence representation. These results underscore that in specialized domains such as healthcare, self-supervised contrastive learning is not optional but decisive, future medical NLP systems should therefore prioritize contrastive frameworks to achieve consistent improvements in classification, clustering, and retrieval.
@@ 4.2 Classification
Our model achieves an F1 score of $0.72\pm 0.23$, outperforming both the strongest non-medical baseline (MPNet Base V2 and Thenlper GTE Base , $0.70\pm 0.22$) and the leading clinical encoder (Abhinand MedEmbed Base and SciNCL, $0.69\pm 0.22$). In practical terms, this +0.02+0.02+0.03+0.03 margin translates to hundreds fewer misclassifications in downstream triage workflows, since each point of F1 corresponds to hundreds of patient records.
@@ 4.3 Clustering
Using the V-measure, our model records $0.38\pm 0.24$, a gain of +0.08+0.08 over the top non-medical contrastive model (MPNet Base V2, $0.30\pm 0.21$) and +0.02+0.02 above the best domain-tuned rival (Abhinand MedEmbed Base , $0.36\pm 0.28$).
@@ 4.4 Pair Classification and Retrieval
For semantically equivalent sentence detection, posts an F1 of $0.74\pm 0.17$, Thenlper GTE Base by +0.05+0.05 and MedEmbed Base by +0.06+0.06. In retrieval, it achieves nDCG@10 $0.45\pm 0.32$, improving over Thenlper GTE Base by +0.04+0.04 and the strongest clinical baseline (Abhinand MedEmbed Base ) by +0.06+0.06. These gains reflect a denser, better-calibrated embedding space that surfaces relevant clinical evidence with fewer query reformulations.
@@ 5.1 Per Source Performance on MedTE
As explained in the data section, we created our benchmark using a diverse set of medical text data from different sources. In table 5, we visualized the average performance of models on different sources of data. Both Abhinand MedEmbed Base and Thenlper GTE Base achieve particularly strong performance on PubMed (0.89–0.90) and on BioRxiv (0.89), substantially outpacing models such as SciBERT (PubMed 0.50; BioRxiv 0.18) and BioClinicalBERT (PubMed 0.47; BioRxiv 0.10). Ours further refines this trend, achieving the highest MIMIC-IV score (0.61) and maintaining robust performance across other clinical sources (Clinical Trials 0.81; MedRxiv 0.74). In comparison, general embeddings such as All MiniLM L6 V2 and MPNet Base V2 hover around 0.51–0.55 on MIMIC-IV and just 0.64–0.67 on Clinical Trials, underscoring the importance of specialized pretraining for domain-specific tasks.
The MedMCQA and MedQA tasks involve multi-option question answering, which poses a significant challenge for embedding models to capture the relationship between a question and its correct answer. All models converge to a 0.39 average on these datasets, highlighting this difficulty. Nevertheless, Ours outperforms competitors on MedQA, likely attributable to its extensive pretraining on a diverse set of medical data. On the general-domain Wikipedia task, domain-adapted models incur only a minor drop, Ours achieves 0.45 and Thenlper GTE Base reaches 0.43, outperforming many general embeddings (All MiniLM L6 V2 at 0.41; BERT Base at 0.26). Overall, these results confirm that our model outperforms existing embeddings across specialized biomedical evaluations while preserving, and in some cases enhancing, performance on broader-scope information sources.
@@
Figure 1: Comparison of model performance on Wikipedia versus MIMIC-IV